OpenStudy (anonymous):

the ratio 18/19 as the sine, cosine or tangent of an angle measure. Find x to the nearest degree.

OpenStudy (anonymous):

So there's to be three values for x, one for each function?

OpenStudy (anonymous):

i.e. sin(x)=18/19, cos(x)=18/19, tan(x)=18/19? Does x even represent an angle here? I don't know what else it could be, but the question could be more clear.

OpenStudy (phi):

You can type into the google search window, for example, asin(18/19) in degrees= or get out your calculator (and make sure it is in "degree mode")